# Approximate Vector Set Search: A Bio-Inspired Approach for High-Dimensional Spaces

### v1

@@ IV-A Overview Algorithm
IV-A1 Binary Hash Encoding for Vector Set
the rationale behind binary hash encoding is rooted in exploiting the inherent similarity-preserving property of hashing to mitigate computational complexity and enhance search efficiency.
First, we define the unified paradigm of LSH shared by various LSH [8, 37, 47, 51] and proposed by [19].
@@ IV-A1 Binary Hash Encoding for Vector Set
Building on this unified paradigm, we adopt a fly-based hashing approach for our framework. As illustrated in Figure 3, this approach fundamentally differs from classical hashing methods in its dimensional transformation strategy. Classical hashing techniques typically perform dimensionality reduction to generate dense hash codes. In contrast, fly-inspired hashing emulates the neural projection patterns in Drosophila’s olfactory circuit [8]. This fly-based approach performs expansion by projections, resulting in sparse binary codes.
The choice of fly-based hashing is motivated by two key factors. First, its sparse binary coding aligns naturally with the set-based structure of Bloom filters, which is essential for our framework’s effectiveness (see Section V). This structural compatibility makes it more suitable than traditional distance-preserving hash functions. Second, among fly-based methods, we specifically adopt BioHash [37] over the original FlyHash [8]. Empirical studies have shown that BioHash achieves approximately twice the similarity preservation performance of FlyHash [37].
@@ 8      $\mathbf{D}^{\mathbf{H}}=\mathbf{D}^{\mathbf{H}}\cup\{\mathbf{H}_{j}\}$;
As illustrated in Figure 4(a), FlyHash [8] generates LSH codes by mimicking the olfactory neural circuit of the fly. Vectors are projected into higher-dimensional spaces using a matrix that simulates projection neurons, enhancing representational capacity. Winner-take-all generates sparse binary codes by selecting the top responding neurons. This has been shown to exhibit locality-sensitive property [8] (Definition 6).
@@
Section IV-B establishes the theoretical foundations of this method. It demonstrates that the hash codes effectively support the approximate top-$k$ vector set search. As illustrated in Figure 4(b), the algorithm quantizes the vector sets into binary hash codes. Next, we introduce the hashing process.
Algorithm 1 outlines the process of generating sparse binary codes for a database of vector sets. The input comprises a database $\mathbf{D}$ of $n$ vector sets $\{\mathbf{V}_{i}\}_{i=1}^{n}$ and a parameter $L_{wta}$ for the winner-takes-all operation, while the output $\mathbf{D}^{\mathbf{H}}$ is a collection of corresponding sparse binary codes. The algorithm iterates through each vector set $\mathbf{V}_{j}$ (line 2), applying the BioHash (see Definition 7) to each vector $\mathbf{v}$ (lines 5-6). This involves a matrix projection $W\mathbf{v}$ followed by the WTA operation, which sets the $L_{wta}$ largest elements to 1 and the rest to 0, resulting in a binary code $\mathbf{h}$ with exactly $L_{wta}$ non-zero elements. These sparse binary codes are aggregated per vector set (line 7) and compiled into $\mathbf{D}^{\mathbf{H}}$ (line 8).
@@ 14       $\mathcal{D}=\mathcal{D}\cup\{(\mathbf{V}_{i},d_{i})\}$;
IV-A2 Search Execution in BioVSS
binary codes offer computational advantages by enabling fast bitwise operations, which are inherently supported by modern CPU architectures [29]. This leads to significant speedups compared to traditional floating-point computations.
Algorithm 2 outlines BioVSS top-$k$ search process in vector set databases. The input includes a query set $\mathbf{Q}$, the number of desired results $k$, the database $\mathbf{D}$, pre-computed sparse binary codes $\mathbf{D}^{\mathbf{H}}$, and the WTA parameter $L_{wta}$. The output $\mathbf{R}$ contains the top-$k$ most similar vector sets. The algorithm begins by generating sparse binary codes for the query set $\mathbf{Q}$ (lines 2-5). This creates the binary representation of the query set $\mathbf{Q}^{\mathbf{H}}$. To leverage CPU-friendly bit operations, we compute the Hamming-based Hausdorff distance $Haus^{H}$ (substituting $dist$ with $hamming(\mathbf{q},\mathbf{v})$ [34] in Definition 4) between $\mathbf{Q}^{\mathbf{H}}$ and each set of binary codes $\mathbf{H}_{j}$ in $\mathbf{D}^{\mathbf{H}}$ (lines 6-8). This step efficiently identifies a set of candidate vector sets $\mathcal{F}$ based on their binary code similarities (line 9). The final stage refines these candidates by computing the actual Hausdorff distance (see Definition 4) between $\mathbf{Q}$ and each candidate vector set $\mathbf{V}_{i}$ in the original space (lines 10-13). The algorithm then selects the top-$k$ results based on these distances (line 14) and then returns $\mathbf{R}$ (line 15).
IV-B Theoretical Analysis of Algorithm Correctness
@@
IV-B1 Bounds on Min-Max Similarity Scores
@@
IV-B2 Upper Tail Probability Bound
@@
IV-B3 Lower Tail Probability Bound
@@
IV-B4 Probabilistic Correctness for Search Results
@@ IV-C Performance Discussion
V Enhancing BioVSS via Cascade Filter
We propose BioVSS++, an enhanced method that addresses the linear scan limitation of BioVSS through a Bio-Inspired Dual-Layer Cascade Filter (BioFilter). The fundamental principle of BioVSS++ exploits the structural similarity between BioHash and Bloom filters, both employing sparse structure for vector encoding and set storage. Next, we detail the filter construction and the search execution.
V-A Filter Construction
The bio-inspired dual-layer cascade filter forms the basis for efficient search execution in BioVSS++. BioFilter consists of two key components: an inverted index based on count Bloom filters and vector set sketches based on binary Bloom filters. The inverted index leverages the count Bloom filter to construct inverted lists, each storing vector sets based on their count values. The vector set sketches employ a binary Bloom filter for second filtering via Hamming distance, enabling efficient scanning. BioFilter’s dual-layer approach significantly enhances search efficiency by reducing the candidate set through successive refinement stages.
@@ V-A Filter Construction
V-A1 Inverted Index Based on Count Bloom Filter
the inverted index serves as the first layer of BioFilter. It accelerates search execution by constructing an inverted index of vector sets to reduce candidate items.
To mitigate the linear scanning limitation of Algorithm 1, we leverage the count Bloom filters to manage vector sets, which form the basis for constructing the inverted index.
@@ 12       Sort $\mathbf{I}_{i}$ in descending order by $c_{i}^{(j)}$;
The inverted index facilitates efficient vector set search. As shown in Figure 5, each vector set is inserted into count Bloom filters. Subsequently, inverted lists are constructed and sorted in descending order based on the count values at various positions in these filters. This approach capitalizes on the principle that higher count values indicate a greater likelihood of collisions at the same positions for similar vector sets.
Algorithm 3 describes the process of generating count Bloom filters for all vector sets. The input consists of a database $\mathbf{D}$ of $n$ vector sets, the winner-takes-all parameter $L_{wta}$, and the length $b$ of the bloom filters. The algorithm begins by generating sparse binary codes for the database using the Gen_Binary_Codes function from Algorithm 1 (line 1). It then constructs count Bloom filters $\mathcal{C}$ for each vector set. Each count Bloom filter $\mathbf{C}^{(j)}$ is created by summing the binary codes of all vectors in the corresponding set $\mathbf{H}_{j}$ (line 2-4) and then returns count Bloom filters (line 5).
Algorithm 4 describes the process of building an inverted index. The input is the same as algorithm 3. The output is an inverted index $\mathbf{I}$. First, count Bloom filters for all data are generated through Algorithm 3 (line 1). For each position $i$ in the count Bloom filters, the algorithm creates an inverted list $\mathbf{I}_{i}$ containing pairs of set indices and their corresponding count values (lines 2-6). Finally, each inverted list is sorted in descending order based on the count values (lines 7-8). This inverted index structure effectively reduces the search space.
V-A2 Vector Set Sketches Based on Binary Bloom Filter
vector set sketches form the second layer of BioFilter, providing a single binary representation of each vector set. These sketches are based on binary Bloom filters, constructed by applying a bitwise OR operation to the binary codes of all vectors within each set.
The binary nature of these sketches enables efficient similarity estimation through Hamming distance, leveraging fast XOR and popcount operations available in modern CPUs [29]. The sketches reduce computational complexity by avoiding the expensive aggregation operations in Hausdorff distance.
@@ V-A2 Vector Set Sketches Based on Binary Bloom Filter
As shown in Figure 5, each vector set is inserted into binary Bloom filters. Subsequently, the binary Bloom filter of each vector set is transformed into a vector set sketch.
@@ 6 $\mathcal{S}=\{\mathbf{S}^{(i)}\}_{i=1}^{n}$, where $\mathbf{S}^{(i)}=\mathbf{B}^{(i)}$;
Algorithm 5 outlines the process of building set sketches in vector set databases. The input consists of a database $\mathbf{D}$ of $n$ vector sets, the winner-takes-all parameter $L_{wta}$, and the length $b$ of the binary code. The output is a collection of set sketches $\mathcal{S}$ that provide compact representations of the vector sets. The algorithm begins by generating sparse binary codes for the database using the Gen_Binary_Codes function from Algorithm 1 (line 1). It then constructs binary Bloom filters $\mathcal{B}$ for each vector set (lines 2-4). Each binary Bloom filter $\mathbf{B}^{(i)}$ is created by applying a bitwise OR operation ($\bigvee$) to all binary codes of vectors in the corresponding set $\mathbf{H}_{i}$. Finally, the algorithm builds the set sketches $\mathcal{S}$ (line 5) by directly using the binary Bloom filters as the sketches and then returns $\mathcal{S}$.
@@ 32       $\mathcal{D}=\mathcal{D}\cup\{(\mathbf{V}_{i},d_{i})\}$;
V-B Search Execution in BioVSS++
@@
This section introduces the search execution process in BioVSS++. Building upon the inverted index and set sketches, BioFilter can quickly filter unrelated vector sets.
@@
The strategy of our search execution lies in a dual-layer filtering mechanism. In the first layer, we employ the inverted index to quickly eliminate a large portion of dissimilar vector sets. In the second layer, we use the vector set sketches for further refinement. This overcomes the limitations of linear scanning in Algorithm 1, reducing the computational overhead.

### v2

@@ 4.1 Overview Algorithm
4.1.1 Binary Hash Encoding for Vector Set
The rationale for employing binary hash encoding lies in harnessing the similarity-preserving nature of hashing to reduce computational overhead and improve search efficiency.
We begin by formalizing the unified framework of LSH, which underpins a range of LSH methods [21, 22, 26, 27], originally introduced in [4].
@@ 4.1.1 Binary Hash Encoding for Vector Set
Building upon this unified paradigm, our framework incorporates a fly-inspired hashing mechanism. As depicted in Figure 4, this method fundamentally contrasts with conventional hashing schemes in its approach to dimensional transformation. While classical hashing methods commonly apply dimensionality reduction to produce dense hash codes, the fly-based strategy draws from the neural projection mechanisms observed in the Drosophila olfactory system [21], performing expansion via projections and yielding sparse binary representations.
The adoption of fly-based hashing is driven by two main considerations. First, its sparse binary encoding naturally aligns with the set-oriented structure of Bloom filters, which is critical for the functionality of our framework (refer to Section 5). This structural alignment renders it more appropriate than traditional distance-preserving hashing functions. Second, within fly-inspired methods, we employ BioHash [22] in preference to the original FlyHash [21], as empirical evaluations indicate that BioHash offers roughly double the similarity preservation performance compared to FlyHash [22].
@@ 8      $\mathbf{D}^{\mathbf{H}}=\mathbf{D}^{\mathbf{H}}\cup\{\mathbf{H}_{j}\}$;
As shown in Figure 5(a), FlyHash [21] constructs LSH codes by simulating the fly’s olfactory neural circuit. It projects input vectors into a higher-dimensional space using a projection matrix that models the behavior of projection neurons, thereby increasing representational richness. Winner-take-all mechanism is then employed to produce sparse binary codes by retaining only the most active neurons. This approach has been proven to satisfy the locality-sensitive property [21] (see Definition 6).
@@
Theoretical justifications for this method are provided in Section 4.2, where we demonstrate that the generated hash codes are well-suited to support approximate top-$k$ vector set search. As depicted in Figure 5(b), the vector sets are first quantized into binary hash codes. We now proceed to describe the hashing procedure in detail.
Algorithm 1 presents the procedure for computing sparse binary representations for a database of vector sets. The input includes a database $\mathbf{D}$ containing $n$ vector sets $\{\mathbf{V}_{i}\}_{i=1}^{n}$ and a parameter $L_{wta}$, which governs the WTA operation. The output $\mathbf{D}^{\mathbf{H}}$ consists of the corresponding sparse binary codes. The algorithm processes each vector set $\mathbf{V}_{j}$ in sequence (line 2), applying the BioHash transformation (refer to Definition 7) to every vector $\mathbf{v}$ it contains (lines 5–6). This transformation involves computing $W\mathbf{v}$ through a linear projection, followed WTA operation that retains the top $L_{wta}$ entries as 1 while setting the rest to 0, yielding a binary vector $\mathbf{h}$ with exactly $L_{wta}$ active bits. These codes are aggregated on a per-vector-set basis (line 7) and stored in $\mathbf{D}^{\mathbf{H}}$ (line 8).
@@ 14       $\mathcal{D}=\mathcal{D}\cup\{(\mathbf{V}_{i},d_{i})\}$;
4.1.2 Search Execution in BioVSS
Binary codes offer computational efficiency by enabling rapid bitwise operations, which are natively supported by modern CPU architectures [44]. This results in substantial acceleration compared to conventional floating-point arithmetic.
Algorithm 2 presents BioVSS procedure for performing top-$k$ search within vector set databases. The inputs include query set $\mathbf{Q}$, target result count $k$, database $\mathbf{D}$, its associated sparse binary representations $\mathbf{D}^{\mathbf{H}}$, and the WTA parameter $L_{wta}$. The output $\mathbf{R}$ comprises top-$k$ vector sets most similar to the query. The algorithm first encodes the query set $\mathbf{Q}$ into its sparse binary form $\mathbf{Q}^{\mathbf{H}}$ (lines 2-5). To take advantage of CPU-optimized bit-level computation, a Hamming-based Hausdorff distance $Haus^{H}$ is computed (using $hamming(\mathbf{q},\mathbf{v})$ [45] in place of $dist$ in Definition 4) between $\mathbf{Q}^{\mathbf{H}}$ and each binary code set $\mathbf{H}_{j}$ in $\mathbf{D}^{\mathbf{H}}$ (lines 6-8). This step yields a candidate set $\mathcal{F}$ of potentially similar vector sets based on binary code proximity (line 9). In the refinement phase, the algorithm evaluates the exact Hausdorff distance (see Definition 4) between $\mathbf{Q}$ and each candidate set $\mathbf{V}_{i}$ in the original vector space (lines 10-13). It then selects the $k$ closest matches (line 14) and returns the final result set $\mathbf{R}$ (line 15).
4.2 Theoretical Analysis of Algorithm Correctness
@@
4.2.1 Bounds on Min-Max Similarity Scores
@@
4.2.2 Upper Tail Probability Bound
@@
4.2.3 Lower Tail Probability Bound
@@
4.2.4 Probabilistic Correctness for Search Results
@@ 4.3 Performance Discussion
5 Enhancing BioVSS via Cascade Filter
We introduce BioVSS++, an improved approach that overcomes the linear scan bottleneck of BioVSS by incorporating a Bio-Inspired Dual-Layer Cascade Filter (BioFilter). The core idea behind BioVSS++ lies in utilizing the structural parallels between BioHash and Bloom filters, both of which rely on sparse representations for encoding vectors and managing set membership. In the following, we elaborate on the construction of the filter and the procedure for executing the search.
5.1 Filter Construction
The bio-inspired dual-layer cascade filter serves as the core mechanism enabling efficient search within BioVSS++. BioFilter is composed of two primary elements: an inverted index built upon count Bloom filters, and vector set sketches constructed using binary Bloom filters. The inverted index utilizes count Bloom filters to generate inverted lists, where vector sets are organized according to their respective count values. For the second-level filtering, the vector set sketches make use of binary Bloom filters, allowing for efficient candidate scanning based on Hamming distance. This two-tiered filtering strategy in BioFilter effectively boosts search performance by progressively narrowing down the candidate pool through sequential refinement.
@@ 5.1 Filter Construction
5.1.1 Inverted Index Based on Count Bloom Filter
The inverted index constitutes the first layer of BioFilter. It improves search efficiency by indexing vector sets in an inverted manner, thereby reducing the number of candidate items.
To overcome the linear scanning bottleneck present in Algorithm 1, we utilize count Bloom filters to manage vector sets, which serve as the foundation for building the inverted index.
@@ 12       Sort $\mathbf{I}_{i}$ in descending order by $c_{i}^{(j)}$;
The inverted index plays a critical role in enabling efficient vector set search. As illustrated in Figure 6, each vector set is first inserted into a count Bloom filter. Inverted lists are then constructed and ranked in descending order according to the count values at each filter position. This strategy leverages the intuition that higher count values imply a greater probability of collisions at the same positions, suggesting stronger similarity among vector sets.
Algorithm 3 outlines the procedure for generating count Bloom filters across all vector sets. The inputs include a database $\mathbf{D}$ containing $n$ vector sets, the winner-takes-all parameter $L_{wta}$, and the desired length $b$ of the Bloom filters. Initially, sparse binary codes for the dataset are generated using the Gen_Binary_Codes procedure defined in Algorithm 1 (line 1). Subsequently, a count Bloom filter $\mathcal{C}$ is computed for each vector set by aggregating the binary codes of its constituent vectors. Specifically, each count Bloom filter $\mathbf{C}^{(j)}$ is produced by summing the binary vectors in $\mathbf{H}_{j}$ (lines 2–4), followed by returning the complete set of filters (line 5).
Algorithm 4 presents the method for constructing the inverted index. This algorithm takes the same input as Algorithm 3, and outputs the inverted index $\mathbf{I}$. First, count Bloom filters for all entries in the dataset are generated using Algorithm 3 (line 1). For each position $i$ in the filters, an inverted list $\mathbf{I}_{i}$ is created, storing pairs of set indices and their corresponding count values (lines 2–6). Finally, each list is sorted in descending order based on these count values (lines 7–8). This index structure effectively narrows down the search space by filtering candidates early in the process.
5.1.2 Vector Set Sketches Based on Binary Bloom Filter
Vector set sketches constitute the second layer of BioFilter, offering a unified binary representation for each vector set. These sketches are generated using binary Bloom filters, where a bitwise OR operation is applied across the binary codes of all vectors within a set.
Due to their binary structure, these sketches allow efficient similarity estimation via Hamming distance, utilizing fast XOR and popcount instructions supported by modern CPU architectures [44]. This design significantly reduces computational overhead by eliminating the need for costly aggregation steps inherent in Hausdorff distance calculations.
@@ 5.1.2 Vector Set Sketches Based on Binary Bloom Filter
As illustrated in Figure 6, each vector set is first inserted into a binary Bloom filter. This binary Bloom filter is then converted into a corresponding vector set sketch.
@@ 6 $\mathcal{S}=\{\mathbf{S}^{(i)}\}_{i=1}^{n}$, where $\mathbf{S}^{(i)}=\mathbf{B}^{(i)}$;
Algorithm 5 details the procedure for constructing set sketches from a vector set database. The input includes a database $\mathbf{D}$ containing $n$ vector sets, the winner-takes-all parameter $L_{wta}$, and the binary code length $b$. The output is a set of sketches $\mathcal{S}$, providing compact encodings for each vector set. The algorithm starts by invoking the Gen_Binary_Codes function from Algorithm 1 to produce sparse binary codes for the entire database (line 1). It then constructs binary Bloom filters $\mathcal{B}$ for all vector sets (lines 2-4), where each binary Bloom filter $\mathbf{B}^{(i)}$ is obtained via a bitwise OR operation ($\bigvee$) over the binary codes of all vectors in the associated set $\mathbf{H}_{i}$. In the final step (line 5), the binary Bloom filters themselves are used directly as vector set sketches $\mathcal{S}$, which are then returned.
@@ 32       $\mathcal{D}=\mathcal{D}\cup\{(\mathbf{V}_{i},d_{i})\}$;
5.2 Search Execution in BioVSS++
@@
This section presents the search execution process in BioVSS++. Leveraging the inverted index and set sketches, BioFilter can efficiently eliminate unrelated vector sets.
@@
Our search execution approach is based on a dual-layer filtering mechanism. At the first layer, we utilize the inverted index to rapidly remove a substantial number of dissimilar vector sets. At the second layer, we apply the vector set sketches for additional filtering. This addresses the drawbacks of linear scanning in Algorithm 1, minimizing the computational overhead.